OpenStudy (anonymous):

Reeti has a mass of 51.0 kg. The Gravitron, a ride that spins so fast that the floor can be removed without the riders falling, reaches a speed of 17.0 m/s and has a radius of 3.20 meters. What is her angular momentum? last q for today/night

hero (hero):

This is so easy

OpenStudy (anonymous):

okay, well can you please help me?

hero (hero):

Yeah, just apply the related formulas

OpenStudy (anonymous):

ok? I'll just figure it out. thanks

hero (hero):

You're welcome

OpenStudy (anonymous):

L = I ω GOT IT

OpenStudy (anonymous):

ω = v/r ω=17/3.20 ω=5.3125 moment of inertia = mass x radius2 I = mr^2 I=51(3.2^2) I=51(10.24) I=522.24 L = I ω L=522.24(5.3125) L=2774.4